# TransVIP: Speech to Speech Translation System with Voice and Isochrony Preservation

### v1

@@ 4.1 Experimental Settings
In this section we introduce the implementation, training end evaluation details of TransVIP system. We conduct our experiment in the English-French mutual translation setting, as it has relatively more data compared to other language pairs..
Implementation Details
All three models within the system are trained using 32 NVIDIA V100 32G GPUs. We utilize Fairseq2 libriary [^3] https://github.com/pytorch/fairseq2, MIT License for model building and the PyTorch Lightning framework [^4] https://lightning.ai/docs/pytorch/stable/, Apache 2.0 License for distributed data parallel (DDP) training. The training time is around one week for each of the three model.
  • Joint Translation model The primary encoder-decoder (Enc-Dec) architecture of the joint model is initialized using the SeamlessM4T S2T model, while the other submodules are trained from scratch. The first-layer RVQ codes are integrated into the vocabulary of foundational S2T model. As mentioned in Section 3.1.3, this model is trained using multiple datasets, including two S2ST datasets: CVSS-T[44] and SeamlessAlign [^5] https://github.com/facebookresearch/seamless_communication/blob/main/docs/m4t/seamless_align_README.md, one internal ST dataset, and one ASR dataset: Common Voice[45] version 15 (English and French subsets). An offline preprocessing step is performed to convert the speech labels into a discrete codec form. During this process, we utilize a VAD tool [^6] https://github.com/snakers4/silero-vad, MIT License to eliminate silence at the beginning and end of the speech recordings. This VAD tool is also used to generate the isochrony embeddings during both training and inference phases.The input to the speech semantic encoder is Fbank features and text is tokenized into BPE, which is consistent with the Seamless framework. The acoustic encoder is a six-layer standard Transformer encoder with a hidden size of 1024. Its input features are extracted from raw waveforms by the SASC encoder, providing richer acoustic information than Fbank features. For decoding we employee beam search algorithm with beam size set to be 5.
  • NAR Acoustic Model This 12-layer transformer model is trained from scratch and utilizes mainly two unsupervised corpora: LibriLight[46], containing 52k hours of English speech, and VoxPopuli[47] French subset, encompassing 23k hours of French speech. We also include audios from SeamlessAlign and Common Voice into training. The datasets are weighted differently during training to ensure that data from both languages are sampled with equal probability. In inference we employee the LBS mentioned in section 3.3. There is 16 RVQ layers in the model, which mean 15 generation steps. The beam size is 10, the sampling number is 20 and $K$ is 3.
  • SASC Model We train the model on audio with a frequency of 16k Hz. The downsample rate is set at 320, resulting in 50 tokens/frames per second, instead of the 75 tokens used in the official DAC checkpoint. This adjustment is primarily made to ensure a consistent token rate that matches the semantic token rate, thereby facilitating the distillation process. The model is trained with in total 24 RVQ codebook while only 16 are employed in the translation system. The SASC model is trained using the full set of the Common Voice version 15 corpus, which includes hours of multilingual speech data. During each training iteration, a random 2-second segment is cropped from the speech samples.
@@
For speech to speech translation translaiton evaluation, we use a subset of CVSS-T [8] fr-en test set containing 300 utterance. It is the first three hundred rows in the original table [^7] https://dl.fbaipublicfiles.com/covost/covost_v2.fr_en.tsv.tar.gz for test set provided by CoVoST 2 [48]. We use source speech as prompt. And all the similarity metrics are also calculated with the source speech as reference. We use at most 10-second prompt for joint translation model and 5-second prompt for NAR acoustic model. If the source speech shorter than that limitation we just use the whole utterance as prompt. For baseline we compared our model with a textless S2UT model and two Seamless models. The textless S2UT model is the All-to-English XM Transformer [^8] https://github.com/facebookresearch/fairseq/tree/ust/examples/speech_matrix release by SpeechMatrix[49]. For seamless models, We compared with two versions, the medium version and the expressive version. The medium version uses the same speech to text translation model as the one we use to initialize our joint translation model, while expressive version has larger model size and stronger text-to-speech model. For codec evaluation, we employee LibriSpeech [50] test-clean, which is widely adopt for codec and TTS evaluation. We compare our SASC with Encodec [21], Speechtokenizer [9] and DAC[22]. We evaluate S2ST performance on translation(BLEU), speaker&prosody similarity(SIM & AutoPCP), Isochrony control(Rate, Pause & SLC) and naturalness. The metrics are described in detail in appendix A.
@@ 4.2 S2ST Evaluation Result
In this section we will show the evaluation result of TransVIP in terms of translation performance, voice preservation, isochrony control and speech naturalness. Ablation study on the design of TransVIP is shown in appendix B.
Table 1: The subjective evaluation results of our model and baseline model. In all scores, higher values are better. The results of baseline models are inferred from official checkpoints. We have also conducted statistical significance test between TransVIP and SeamlessExpressive at 0.05 significance level. Any results that are statistically significantly better are highlighted in bold. Abbreviation: A.PCP(AutoPCP), Nat.(Natureness)
@@
We evaluated the translation performance using BLEU and ASR-BLEU metrics as presented in Table 1. 1) By comparing our models, TransVIP and the Seamless baseline, with the state-of-the-art (SOTA) textless S2UT model and analyzing the generated samples, we observed that both our models significantly outperform the textless S2UT model. Notably, the textless model exhibits severe issues with repetition and "illusion" (generating audio that seems plausible but is unrelated to the input), which have been substantially mitigated in both the Seamless model and our TransVIP. This confirms our hypothesis that the inclusion of an intermediate text is crucial for successful S2ST generation. 2) Our TransVIP demonstrates highly competitive results compared to the Seamless baselines. In the French-to-English translation direction, our model achieved a BLEU score of 32.60, which is 3.65 points higher than the SeamlessM4T medium and 1.75 points above the larger SeamlessExpressive. In the English-to-French direction, our model surpasses the SeamlessM4T medium by 6.22 points. Additionally, the ASR-BLEU performance of our model is comparable to that of SeamlessExpressive, with only a slight difference of -0.11 points despite a 1.87-point gap in the text BLEU score. The disparity in text BLEU scores is attributed to differences in model size, and our limited S2TT data precludes rectifying this issue.
@@
We assess speaker similarity in the SIM column of Table 1. The Textless S2UT model and SeamlessM4T Medium do not effectively preserve speaker identity, as evidenced by their near-zero similarity scores. In contrast, our TransVIP model achieves scores of 0.320 and 0.395 in the fr-en and en-fr directions, respectively, slightly outperforming the SeamlessExpressive baseline, which scores 0.256 and 0.335. Notably, TransVIP also exceeds the fr-en target speech from the CVSS-T dataset, which is used in our training, by 0.115 points. This indicates that our model’s similarity performance transcends the limitations of its training dataset. Regarding prosody similarity, TransVIP’s AutoPCP score surpasses both Textless S2UT and SeamlessM4T Medium. Compared to SeamlessExpressive, TransVIP performs 0.25 points lower in the fr-en direction but 0.14 points higher in en-fr. It is significant to note that unlike TransVIP, SeamlessExpressive explicitly incorporates a prosody encoder.
@@
We observed that the quality of the conditioning speech prompt significantly influences the naturalness of the generated speech, as measured by the NISQA score. For instance, when processing low-quality real French data (Fr-En), SeamlessExpressive registers the lowest score at 2.91. Conversely, with high-quality synthesized English data (En-Fr) as input, SeamlessExpressive achieves the highest score of 3.57. Meanwhile, our TransVIP model demonstrates consistent performance, scoring 3.19 in the Fr-En direction and 3.40 in the en-fr direction. Given that the NISQA score can present difficulties in evaluating the quality or naturalness of noisy speech, we also assess the naturalness of speech using a subjective metric known as the Comparative Mean Opinion Score (CMOS). For this evaluation, 12 professional linguistic experts judged a total of 120 pairs of utterances translated by TransVIP and SeamlessExpressive. The results indicated that TransVIP’s performance is significantly better (p<0.01) than that of SeamlessExpressive, i.e., TransVIP outperforms SeamlessExpressive by achieving a preference score that is higher by 0.46. The instruction of subjective test is shown in Appendix F.
@@ 4.3 Codec Evaluation Result
We show the evaluation result in table 4 and group the result by bandwidth(BW). The SpeechTokenizer that has underwent semantic distillation perform well in NISQA score, showing that it can produce a clean and nature voice. On the other hand, DAC perform well in similarity and WER, showing a strong capability of a faithful reconstruct of original speech. Finally, our SASC successfully combines the strength the two codec, showing a very competitive result in all of the four metrics.
Table 2: Codec re-synthesize evaluation result. The codecs are grouped by bandwidth. * means the result is infered from official checkpoint. †means it is reproduced result using official recipe and Common Voice data. Abbreviation: NQ(number of quantizer), FR(frame rate), BW(Bandwidth), Nat.(Natureness)
@@
5 Conclusion
In this paper, we develop a speech-to-speech translation framework that 1) can preserve speaker voice and isochrony to generate high quality translated speech, which is practical in both daily communication and automatic video dubbing. 2) can inference in an end to end manner, which jointly consider the text and the speech probability during the inference phase and make fully use various data during the training phase. In the Fr-En direction TransVIP surpasses the current SOTA model with comparable translation performance and much improved voice and isochrony preservation.
Limitations and future works
1) Performance & Language support: for the resource and data limited currently we only trained our model in Fr-En language pair. Currently we only use around 5k hours audio to train the translation model compare to more than 50k in previous works[7, 40, 38], which indicates a potential performance improvement if we scale up the data. Also, In the future we hope to extend our framework to many-to-many setting by large-scale multi-lingual training, just like seamless does. 2) More detailed attribute control: On reviewing the test sample we find sometimes TransVIP mistakenly change the tone of the speech, like translating interrogative sentences into declarative sentences. It is reasonable to hypnosis that a more detailed control on more attributes like intonation or emotion could help to solve the problem.
Acknowledgments and Disclosure of Funding
We extend our gratitude to Chung-Hsien Tsai, Canrun Li, and Zhen Xiao for their invaluable contribution in conducting the subjective evaluation and for their efforts in assembling the video dubbing material. Also we appreciate Microsoft Azure for providing computing services and GPUs.
Appendix A Evaluation Metrics
For S2ST evaluation, we evaluate translation performance (BLEU and ASR-BLEU), speaker similarity (SIM), prosody similarity(AutoPCP), isochrony control (Rate, Pause and SLC-0.2/0.4) and naturalness (NISQA). The score the higher the better for all the S2ST metrics.
  • BLEU is calculated between text output and text label using the ’corpus_bleu’ method of sacrebleu [^9] https://github.com/mjpost/sacrebleu, Apache 2.0 License.[51]. To calculate speech BLEU(also known as ASR-BLEU), we utilize a whisper large v3 model to transcribe the generated speech into text then calculate BLEU score between transcribed text and text label.
  • For naturalness evaluation we use NISQA-TTS(V1.0) [^12] https://github.com/gabrielmittag/NISQA, MIT License, which is a model to evaluate the naturalness and quality of synthesized speech. [55]
For codec evaluation, we use metrics that are commonly employed in zero shot-TTS to evaluate the re-synthesize quality of the codec, which are 1) SIM, 2) NISQA, 3) WER. The NISQA metrics is the same as those in above S2ST evaluation. A market-leading Speech Recognition API is used for Word Error Rate (WER) calculation. It measures the correctness and intelligibility for re-synthesized speech. We use two similarity scores. SIM is used to calculate the similarity between original speech and resynthesised speech. Generally, this SIM value is very high because the two embeddings used in the calculation are extracted from speech that shares the same content in the same language. Following [10, 56], the SIM-o metric is computed by first segmenting the audio into two parts: a 3-second prompt and the remaining portion. The latter part is then re-synthesized. Finally, similairy score is calculated between the re-synthesized audio and the original prompt.
Appendix B Ablation Study
B.1 Ablation on Acoustic Embedding
In the process of semantic distillation within the codec, a significant amount of acoustic information was disentangled from the first-layer codec. This leads to the question: Is the acoustic encoder still necessary in the joint translation model? To investigate this, we leveraged the model’s capability to perform inference without the acoustic embedding. We conducted such inferences and compared the outcomes, as shown in 3. Without the acoustic embedding, speaker similarity scores decreased by 0.040 in the French-English (Fr-En) translations and by 0.032 in the English-French (En-Fr) translations. Additionally, there was a slight decline in the AutoPCP scores. These findings suggest that residual acoustic information remains within the first-layer codec, underscoring the necessity of including a prompt embedding.
@@ B.1 Ablation on Acoustic Embedding
B.2 Ablation on the Choice of Codec

### v2

@@ 5.1 Experimental Settings
In this section, we introduce the implementation, training, and evaluation of the TransVIP system. We conducted our experiment in the English-French mutual translation setting, as it had relatively more data than other language pairs. Appendix A and B include more details about our models and metrics.
Implementation
All three models within the system are trained using 32 NVIDIA V100 32G GPUs. We utilize Fairseq2 libriary [^3] https://github.com/pytorch/fairseq2, MIT License for model building and the PyTorch Lightning framework [^4] https://lightning.ai/docs/pytorch/stable/, Apache 2.0 License for distributed data parallel (DDP) training. The training time for each of the three models is around one week.
@@
For speech-to-speech translation translation evaluation, we use a subset of CVSS-T [8] fr-en test set containing 300 utterances. It is the first three hundred rows in the original table [^5] https://dl.fbaipublicfiles.com/covost/covost_v2.fr_en.tsv.tar.gz for test set provided by CoVoST 2 [48]. We use source speech as the prompt. All the similarity metrics are also calculated with the source speech as the reference.
We use at most, a 10-second prompt for the joint translation model and a 5-second prompt for the NAR acoustic model. If the source speech is shorter than that limitation we just use the whole utterance as the prompt. For baseline, we compared our model with a textless S2UT model and two Seamless models. The textless S2UT model is the All-to-English XM Transformer [^6] https://github.com/facebookresearch/fairseq/tree/ust/examples/speech_matrix release by SpeechMatrix[49]. For seamless models, We compared with two versions, the medium version, and the expressive version. The medium version uses the same speech-to-text translation model as the one we use to initialize our joint translation model, while the expressive version has larger model size and a stronger text-to-speech model. For codec evaluation, we employ LibriSpeech [50] test-clean, which is widely adopted for codec and TTS evaluation. We compare our SASC with Encodec [21], Speechtokenizer [9] and DAC[22]. We evaluate S2ST performance on translation(BLEU), speaker&prosody similarity(SIM & AutoPCP), Isochrony control(Rate, Pause & SLC), and naturalness.
@@ 5.2 S2ST Evaluation Result
In this section, we will show the evaluation result of TransVIP in translation performance, voice preservation, isochrony control, and speech naturalness.
Table 1: The subjective evaluation results of our model and baseline model. In all scores, higher values are better. The results of baseline models are inferred from official checkpoints. We have also conducted a statistical significance test between TransVIP and SeamlessExpressive at a 0.05 significance level. Any results that are statistically significantly better are highlighted in bold. Abbreviation: A.PCP(AutoPCP), Nat.(Natureness)
@@
We evaluated the translation performance using BLEU and ASR-BLEU metrics as presented in Table 1. 1) By comparing our models, TransVIP and the Seamless baseline, with the state-of-the-art (SOTA) textless S2UT model and analyzing the generated samples, we observed that both our models significantly outperform the textless S2UT model. Notably, the textless model exhibits severe issues with repetition and "illusion" (generating audio that seems plausible but is unrelated to the input), which have been substantially mitigated in both the Seamless model and our TransVIP. This confirms our hypothesis that including an intermediate text is crucial for successful S2ST generation. 2) Our TransVIP demonstrates highly competitive results compared to the Seamless baselines. In the French-to-English translation direction, our model achieved a BLEU score of 32.60, 3.65 points higher than the SeamlessM4T medium and 1.75 points above the larger SeamlessExpressive. In the English-to-French direction, our model surpasses the SeamlessM4T medium by 6.22 points. Additionally, the ASR-BLEU performance of our model is comparable to that of SeamlessExpressive, with only a slight difference of -0.11 points despite a 1.87-point gap in the text BLEU score. The disparity in text BLEU scores is attributed to differences in model size, and our limited S2TT data precludes rectifying this issue.
@@
We assess speaker similarity in the SIM column of Table 1. The Textless S2UT model and SeamlessM4T Medium do not effectively preserve speaker identity, as evidenced by their near-zero similarity scores. In contrast, our TransVIP model achieves scores of 0.320 and 0.395 in the fr-en and en-fr directions, respectively, slightly outperforming the SeamlessExpressive baseline, which scores 0.256 and 0.335. TransVIP also exceeds the fr-en target speech from the CVSS-T dataset, used in our training, by 0.115 points. This indicates that our model’s similarity performance transcends the limitations of its training dataset. Regarding prosody similarity, TransVIP’s AutoPCP score surpasses Textless S2UT and SeamlessM4T Medium. Compared to SeamlessExpressive, TransVIP performs 0.25 points lower in the fr-en direction but 0.14 points higher in en-fr. It is significant to note that, unlike TransVIP, SeamlessExpressive explicitly incorporates a prosody encoder.
@@
We observed that the quality of the conditioning speech prompt significantly influences the naturalness of the generated speech, as measured by the NISQA score. For instance, when processing low-quality real French data (Fr-En), SeamlessExpressive registers the lowest score at 2.91. Conversely, with high-quality synthesized English data (En-Fr) as input, SeamlessExpressive achieves the highest score of 3.57. Meanwhile, our TransVIP model demonstrates consistent performance, scoring 3.19 in the Fr-En direction and 3.40 in the en-fr direction. Given that the NISQA score can present difficulties in evaluating the quality or naturalness of noisy speech, we also assess the naturalness of speech using a subjective metric known as the Comparative Mean Opinion Score (CMOS). For this evaluation, 12 professional linguistic experts judged 120 pairs of utterances translated by TransVIP and SeamlessExpressive. The results indicated that TransVIP’s performance is significantly better (p<0.01) than SeamlessExpressive, i.e., TransVIP outperforms SeamlessExpressive by achieving a preference score higher by 0.46. The instruction for the subjective test is shown in Appendix G.
@@ 5.3 Codec Evaluation Result
We show the evaluation result in Table 6 and group the result by bandwidth(BW). The SpeechTokenizer that has undergone semantic distillation performs well in the NISQA score, producing a clean and natural voice. On the other hand, DAC performs well in similarity and WER, showing a strong capability of a faithful reconstruction of the original speech. Finally, our SASC successfully combines the strength of two codecs, showing a very competitive result in all four metrics.
Table 2: Codec re-synthesize evaluation result. The codecs are grouped by bandwidth. * means the result is inferred from the official checkpoint. †means it is reproduced result using the official recipe. Abbreviation: NQ(number of quantizers), FR(frame rate), BW(Bandwidth), Nat.(Natureness)
@@
6 Ablation Study
6.1 Ablation on Acoustic Embedding
In the process of semantic distillation within the codec, a significant amount of acoustic information was disentangled from the first-layer codec. This raises the question of whether the acoustic encoder is still necessary in the joint translation model. To investigate this, we leveraged the model’s capability to perform inference without the acoustic embedding. We conducted such inferences and compared the outcomes, as shown in Table 3. Without the acoustic embedding, speaker similarity scores decreased by 0.040 in the French-English (Fr-En) translations and by 0.032 in the English-French (En-Fr) translations. Additionally, there was a slight decline in the AutoPCP scores. These findings suggest that residual acoustic information remains within the first-layer codec, underscoring the necessity of including a prompt embedding.
@@ 6.1 Ablation on Acoustic Embedding
6.2 Ablation on Isochrony Control method
We compared our proposed Isochrony control method with and without using other strategies. We presented the results in Table 4, where a) No Isochrony Control (No IC). b) Isochrony Control on the Decoder (Dec IC). This involves adding the Isochrony embedding to the input of the encoder as another positional embedding. We implemented the method from [54] in our system. c) Isochrony Control on the Decoder with Future Pause Information (Dec IC + FPI). This is an improvement over (b). In addition to the distance to the global end and VAD information, two extra pieces of information are encoded: the distance to the next pause and the number of pauses in the future. We implemented the method from [pal2023improving] in our system.
It demonstrates that our proposed method reaches the best isochronic control. The results also show that this approach can improve the ASR-BLEU score compared to Isochrony control on the decoder, meaning the model is more confident and accurate in the generation and makes fewer errors like repetition and truncation.
Table 4: Ablation study on the isochrony control strategy.
ASR-BLEU Overlap SLCp
0.2 0.4
No IC 30.81 0.689 0.63 0.87
Dec IC 30.51 0.748 0.75 0.90
Dec IC+ FPI 30.45 0.766 0.77 0.91
Enc IC (Proposed) 30.62 0.784 0.82 0.95
6.3 Ablation Study on the NAR Acoustic Model
The textless model is trained on a combination of datasets including Librilight, VoxPopuli French subset, SeamlessAlign, and Common Voice English and French subsets. As some datasets do not provide text transcriptions, we substitute the LibriHeavy dataset in the model trained with text input. Additionally, in the textless model, long audio segments are randomly clipped, whereas in the model with text input, clips strictly adhere to timestamps provided in the metadata to align with text transcriptions. We compare the Fr-En uni-directional setup to mitigate the impact of the absence of a large-scale French dataset on the model with text input. Despite the apparent data discrepancy, textless modeling demonstrates superior performance, underscoring one of its key data accessibility advantages. All other hyperparameters, including training steps, remain consistent across both models.
NAR Model Datasets SIM ASR-BLEU Nat.
NAR w/o text LL,VP,SA,CV 0.320 32.60 3.19
   - LBS 0.309 32.30 3.17
NAR w/ text LH,SA,CV 0.307 31.52 3.10
   - LBS 0.298 31.03 3.09
7 Conclusion, Limitations and future works
In this paper, we develop a speech-to-speech translation framework that preserves speaker voice and isochrony, generating high-quality translated speech suitable for both daily communication and automatic video dubbing. Our framework can infer in an end-to-end manner, jointly considering text and speech probability during the inference phase and making full use of various data during the training phase. In the Fr-En direction, TransVIP surpasses the current state-of-the-art model with comparable translation performance and significantly improved voice and isochrony preservation.
Limitations and future works: 1) Performance & Language Support: Due to current resource and data limitations, we have only trained our model on the Fr-En language pair. We used approximately 5k hours of audio to train the translation model, compared to over 50k hours in previous works[7, 40, 38]. This suggests that there is potential for performance improvement if we scale up the data. Additionally, we aim to extend our framework to many-to-many settings through large-scale multilingual training, similar to Seamless. 2) More Detailed Attribute Control: Upon reviewing the test samples, we found that TransVIP occasionally alters the tone of the speech, such as translating interrogative sentences into declarative ones. It is reasonable to hypothesize that more detailed control of attributes like intonation or emotion could help address this issue.
Broader Impact: TransVIP was developed to improve cross-lingual communication and assist with automated video dubbing workflows. However, it carries potential risks, including the misuse of the model for impersonating specific speakers by inputting targeted text and speaker prompts. Additionally, as a generative model, there is a possibility it may produce toxic or biased outputs, although such occurrences were not observed in our test samples. It is imperative to monitor for any misuse or harmful outputs rigorously. We must also allow users to report such incidents, ensuring responsible usage and continuous system improvement.
Acknowledgments and Disclosure of Funding
We extend our gratitude to Chung-Hsien Tsai, Canrun Li, and Zhen Xiao for their invaluable contribution in conducting the subjective evaluation and for their efforts in assembling the video dubbing material. Also, we appreciate Microsoft Azure for providing computing services and GPUs.
Appendix A Evaluation Metrics
For S2ST evaluation, we evaluate translation performance (BLEU and ASR-BLEU), speaker similarity (SIM), prosody similarity(AutoPCP), isochrony control (Rate, Pause, and SLC-0.2/0.4), and naturalness (NISQA). The score is the higher the better for all the S2ST metrics.
  • BLEU is calculated between text output and text label using the ’corpus_bleu’ method of sacrebleu [^7] https://github.com/mjpost/sacrebleu, Apache 2.0 License.[51]. To calculate speech BLEU(also known as ASR-BLEU), we utilize a whisper large v3 model to transcribe the generated speech into text and then calculate the BLEU score between the transcribed text and text label.
  • For naturalness evaluation, we use NISQA-TTS(V1.0) [^10] https://github.com/gabrielmittag/NISQA, MIT License, which is a model to evaluate the naturalness and quality of synthesized speech. [55]
For codec evaluation, we use metrics that are commonly employed in zero shot-TTS to evaluate the re-synthesize quality of the codec, which are 1) SIM, 2) NISQA, and 3) WER. The NISQA metrics are the same as those in the above S2ST evaluation. A market-leading Speech Recognition API is used for Word Error Rate (WER) calculation. It measures the correctness and intelligibility of re-synthesized speech. We use two similarity scores. SIM is used to calculate the similarity between original speech and resynthesised speech. Generally, this SIM value is very high because the two embeddings used in the calculation are extracted from speech that shares the same content in the same language. Following [10, 56], the SIM-o metric is computed by first segmenting the audio into two parts: a 3-second prompt and the remaining portion. The latter part is then re-synthesized. Finally, similairy score is calculated between the re-synthesized audio and the original prompt.
Appendix B Implementaion Details
Joint Translation model
The primary encoder-decoder (Enc-Dec) architecture of the joint model is initialized using the SeamlessM4T S2T model, while the other submodules are trained from scratch. The first-layer RVQ codes are integrated into the vocabulary of the foundational S2T model. As mentioned in Section 3.4, this model is trained using multiple datasets, including two S2ST datasets: CVSS-T[44] and SeamlessAlign [^11] https://github.com/facebookresearch/seamless_communication/blob/main/docs/m4t/seamless_align_README.md, one internal ST dataset, and one ASR dataset: Common Voice[45] version 15 (English and French subsets). An offline preprocessing step is performed to convert the speech labels into a discrete codec form. During this process, we utilize a VAD tool [^12] https://github.com/snakers4/silero-vad, MIT License to eliminate silence at the beginning and end of the speech recordings. This VAD tool is also used to generate the isochrony embeddings during both the training and inference phases. The input to the speech semantic encoder is Fbank features and text is tokenized into BPE, which is consistent with the Seamless framework. The acoustic encoder is a six-layer standard Transformer encoder with a hidden size of 1024. Its input features are extracted from raw waveforms by the SASC encoder, providing richer acoustic information than Fbank features. For decoding we employed a beam search algorithm with a beam size set to 5.
NAR Acoustic Model
This 12-layer transformer model is trained from scratch and utilizes mainly two unsupervised corpora: LibriLight[46], containing 52k hours of English speech, and VoxPopuli[47] French subset, encompassing 23k hours of French speech. We also included audio from SeamlessAlign and Common Voice in the training. The datasets are weighted differently during training to ensure that data from both languages are sampled with equal probability. In inference, we employ the LBS mentioned in section 4.2. There are 16 RVQ layers in the model, which means 15 generation steps. The beam size is 10, the sampling number is 20 and $K$ is 3.
SASC Model
We train the model on audio with a frequency of 16k Hz. The downsample rate is set at 320, resulting in 50 tokens/frames per second, instead of the 75 tokens used in the official DAC checkpoint. This adjustment is primarily made to ensure a consistent token rate that matches the semantic token rate, thereby facilitating the distillation process. The model is trained with 24 RVQ codebooks while only 16 are employed in the translation system. The SASC model is trained using the full set of the Common Voice version 15 corpus, which includes hours of multilingual speech data. During each training iteration, a random 2-second segment is cropped from the speech samples.
Appendix C Additional Ablation on the Choice of Codec